If the distance between the sun and the earth is increased by three times then the attraction between the two will

Grade:11

1 Answers

Arun
6 years ago

The formula governing gravtational attraction between two bodies, each with uniform spherical distribution of mass (which is approximately the case for the sun and the earth) is: F = [-G(Mm/r^2)]

Now f r is increased by three times, it becomes 3r and the new value of force will be [-G{Mm/(3r)^2}] = [-G(Mm/r^2)]/9 = F/9.

Thus, the attractive force becomes a nineth of its former value.